OpenStudy (anonymous):

can someone plz explain and help me put this into standard form >>> Y=0.05x^2+x+1?

OpenStudy (anonymous):

Y=0.05x^2+x+1 This is in standard form for a parabola, which you have. Do you mean vertex form? vertex form = y= a(x-h)2+k

OpenStudy (anonymous):

yes thank you now I know why no one answers my question srry.

OpenStudy (anonymous):

Ok one sec let me type it out. You just want the answer or would you want me to explain?

OpenStudy (anonymous):

explain. and show.

OpenStudy (anonymous):

kk give me a sec

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

is the vertex (.25,0)?

OpenStudy (anonymous):

Grrr was typing out a huge tutorial and some how it got wiped lol

OpenStudy (anonymous):

Got to the x of the vertex, which will be 10

OpenStudy (anonymous):

omg. that sucks.srry

OpenStudy (anonymous):

so my answer is wrong?

OpenStudy (anonymous):

Ok we need to find the vertex. To find the vertex we use the Y=0.05x^2+x+1 equation and the formula \( \frac{b}{2(a)} \) b = x =1 a = .05 \[ \frac{1}{2(.05)} \] \[ \frac{1}{.1} = 10 \] Ok our x for our vertex = 10 now to find the y for our vertex point we need to insert 10 for x in the equation Y=0.05x^2+x+1 \[ Y=0.05(10)^2+10+1 \] \[ Y=0.05(100)+10+1 \] \[ Y=5+10+1 \] \[ Y=11 \] Now we have our y = 11 so our vertex is (10,11) so to put this in vertex form we plug our points in to y= a(x-h)^2+k a = 0.05 x = x h = vertex x 10 k = vertex y 11 Now just plugin \[y= a(x-h)^2+k \] \[y= 0.05(x-10)^2+11 \]

OpenStudy (anonymous):

just give me a sec to read it over plz

OpenStudy (anonymous):

Sorry, where ever you see 11 it should be 16

OpenStudy (anonymous):

I was rushing that time around lol

OpenStudy (anonymous):

oh okay so y=16 then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol that's fine , I'm glad you took the time to help me out.

OpenStudy (anonymous):

so the axis is -16?

OpenStudy (anonymous):

Also, remember the x of the vertex can be found using \( \frac{b}{2(a)} \) which will give you the x of the vertex

OpenStudy (anonymous):

no wait the axis would be -10.

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank god.

OpenStudy (anonymous):

okay , well thanks so much for helping me.

OpenStudy (anonymous):

No problem anytime.

OpenStudy (anonymous):

I am here to help with what I can

OpenStudy (anonymous):

I know math can be frustrating lol

OpenStudy (anonymous):

yah it really is , I have a short fuse for this stuff sometimes. lol

OpenStudy (anonymous):

It is a lot harder too and more frustrating when you don't have anyone around to ask questions. That is why I like places like this :-)

OpenStudy (anonymous):

ikr , I totally agree, I'm glad I found this place.

OpenStudy (anonymous):

Ok have a good one. I will see ya around

OpenStudy (anonymous):

kk. u2.